OpenStudy (anonymous):

Simplify the expression: (x^m)^n × (x^n)^n-m Please explain the steps to doing this!

OpenStudy (jack1):

\[(x^m)^n = x ^{m \times n}\]

OpenStudy (jack1):

\[(x^n)^{(n-m)} = x ^{(n \times n - n \times m)}\]

OpenStudy (jack1):

so x to the power of negative something = 1/x to the power of something example: x^(-2) = 1/(x^2)

OpenStudy (jack1):

so split the above : x^(nn - nm) = x^nn times x^-nm = x^nn / x^nm following ok so far @kelly.dinh1 ?

OpenStudy (jack1):

... u still there kelly?

OpenStudy (anonymous):

Yes, I am. I'm so sorry! Was busy working on another equation.

OpenStudy (anonymous):

I don't understand. Where did the (x^n)(n-m) = x^(n-n - n×m) come from?

OpenStudy (jack1):

these are basic exponent principals (maths rules) go here: http://www.mathsisfun.com/algebra/exponent-laws.html scroll down to the "laws of exponents" section and learn these maths laws u need to understand them before you can understand this type of equation main ones are x^-6 = 1 / x^6 (x^2)^4 = x^(2 times 4) x^2 times x^4 = x^(2+4)

OpenStudy (anonymous):

Alright, I will. It's been awhile since I've touched up on this subject so this review was needed. Thank you so much!

OpenStudy (jack1):

sáll good, happy to help hey slaters